# SMReferences

### v1

@@ Abstract
Finally, a noteworthy feature of our system is that, due to the oblate geometry of the sample, the turbulent flow exhibits mainly two-dimensional characteristics, whereas the mean flow of the collective modes is three-dimensional. This disparity could limit the use of the RANS method, which presupposes three-dimensional homogeneous isotropic turbulence. Previous research on classical turbulence driven by two-dimensional forcing demonstrated that enhancing the three-dimensional nature of turbulent flows can increase the eddy viscosity by suppressing the inverse energy cascade [58]. This suggests that probing the dimensional crossover presents an exciting direction for further investigation.
@@
In summary, we have observed an enhancement in the damping of collective oscillations of BECs in the presence of turbulence. Investigating the viscous effects induced by turbulence through its coupling to collective excitations provides a novel framework for understanding the energy-transfer mechanisms and the hydrodynamic properties of superfluid turbulence. This work can extend to various collective excitation modes, particularly including a persistent monopole mode observed in spherically symmetric three-dimensional BECs [59].
@@ Acknowledgements.
This work was supported by the National Research Foundation of Korea (Grants No. NRF-2023R1A2C3006565, No. RS-2023-NR119928, and No. RS-2024-00413957).
Supplemental Material
.1 Methods
.1.1 Sample preparation
We prepared a thermal cloud of 23Na atoms in the $|F=1,m_{F}=-1\rangle$ state in an optical dipole trap (ODT) and applied evaporative cooling by reducing the trap depth to produce a Bose-Einstein condensate (BEC). The thermal fraction $\zeta_{\text{th}}$ of the sample was controlled by adjusting the final value $U_{\text{f}}$ of the trap depth. At our lowest temperatures, the typical number of atoms for the BEC was approximately $7\times 10^{6}$. The atomic sample was held in the ODT for several seconds to relax residual motion before driving the collective excitations of the trapped BEC.
Depending on the target experiment, the atomic sample was transformed into a two-component system with $|m_{F}=\pm 1\rangle$ states ($S=2$) or a turbulent system with $|m_{F}=0,\pm 1\rangle$ states ($S=3$). For the two-component system, first we transferred atoms to the $|m_{F}=0\rangle$ state using a rapid adiabatic passage with an RF magnetic field, with this followed by a pulse of a strong magnetic field gradient to purify the spin state. We then applied a $\pi/2$ resonant RF magnetic pulse to form an equal mixture of the $m_{F}=\pm 1$ components. After the preparation of the spin mixture, we irradiated a continuous microwave field detuned from the $F=1\rightarrow F=2$ transition to shift the quadratic Zeeman energy to a small negative value. This prevents the creation of atoms in the $|F=1,m_{F}=0\rangle$ state via spin exchange processes, stabilizing the two-component system [1, 2].
The turbulent sample was prepared with the continuous application of the resonant RF field [3]. A steady turbulent state with an irregular spin texture was generated within 0.5 s, exhibiting equal populations of the three $m_{F}=0,\pm 1$ spin components. This turbulence is sustained by chaotic spin dynamics under RF spin driving [4], as detailed in Section B. Following 2 s of RF driving, we conducted the collective oscillation experiment while maintaining the RF spin driving. The BEC lifetime exceeded 30 s under RF driving ($S=3$) and was approximately 5 s under the microwave dressing ($S=2$), significantly longer than our measurement duration, which lasted for a few hundreds of ms.
@@
.1.2 Imaging
An absorption image of the sample was taken in the $z$ direction after a time of flight of $\tau=18$ ms by releasing the trapping potential. In the case shown in Fig. 3(a), a magnetic field gradient pulse was applied during the expansion to separate the different spin components spatially. The number of atoms contained in the thermal gas, $N_{\text{th}}$, was estimated from a two-dimensional Gaussian distribution fit to the outer region of the cloud, and the number of condensed atoms $N_{c}$ was determined from an image obtained by subtracting the fitted thermal profile from the original image. The thermal fraction is given by $\zeta_{\text{th}}=N_{\text{th}}/N$ with $N=N_{c}+N_{\text{th}}$. The width $W_{\alpha}(t)$ of the condensates along the $\alpha$ axis ($\alpha=x,y,z$) was determined by fitting the two-dimensional inverted parabola to the subtracted profile.
.1.3 Analysis of width oscillations
@@
The second term accounts for the width change during the time of flight, owing to the oscillation velocity of the condensate. Here we neglect the expansion effect due to the mean-field energy, which is negligibly small along the $y$ direction for our highly oblate geometry. When a trapped BEC oscillates as $W_{i}(t)=W_{\text{eq}}(1+A\sin(\omega t+\phi))$, using Eq. (S1), the measured width is given by $W(t)=W_{\text{eq}}(1+\alpha_{\tau}A\sin(\omega t+\phi+\varphi_{\tau}))$ with $\alpha_{\tau}=\sqrt{1+(\omega\tau)^{2}}$ and $\varphi_{\tau}=\tan^{-1}(\omega\tau)$. $W_{\text{eq}}$ denotes the in-situ width of the condensate at equilibrium.
In our data analysis, we used the normalized width $\tilde{W}=W/W_{\text{eq}}$ to account for variations in the number of atoms in the sample. Using the relationship of $W_{\text{eq}}\propto U_{\text{f}}^{-1/5}N_{c}^{1/5}$ in the Thomas-Fermi (TF) approximation, the equilibrium width for a given $N_{c}$ was estimated as $W_{\text{eq}}(N_{c})=\overline{W}_{\text{eq}}(N_{c}/\overline{N_{c}})^{1/5}$, where $\overline{W}_{\text{eq}}$ and $\overline{N_{c}}$ are the mean width and the mean number of atoms of the condensate, as determined by averaging tens of measurements without trap modulations.
.2 Stationary spin-superflow turbulence
@@
×$\bm{\nabla}\times\bm{v}=-\frac{i\hbar}{m}\bm{\nabla}\bm{\zeta}^{\dagger}\times\bm{\nabla}\bm{\zeta},$ (S3)
@@
.3 Quadrupole modes of highly oblate condensates
@@
Figure S1: (a) Quadrupole oscillation modes for an oblate BEC with trapping frequencies $\omega_{x}<\omega_{y}<\omega_{z}$. The oscillatory motion is characterized by $(b_{x},b_{y},b_{z})$, as shown in Eq. (S9), and the three modes are denoted by X, Y, and Z, respectively, corresponding to the dominant oscillation axis. The variation in the density profile in the $xy$ and $xz$ planes for each mode is presented. (b) and (c) Experimental observation of the quadrupole oscillations of BECs. BEC samples with $S=1$ were prepared at our lowest temperatures and the X or Y mode was selectively excited by short trap modulations at a frequency slightly red-detuned from resonance. The upper panels in (b) and (c) display time-of-flight images of BECs oscillating in the X and Y modes, respectively, at different hold times. The lower panels show the time evolution of the normalized condensate widths in the $x$ and $y$ directions, respectively. Solid lines indicate damped sinusoidal fits to the experimental data.
@@
The three quadrupole excitation modes for an oblate trapping potential with $\omega_{x}<\omega_{y}<\omega_{z}$ are depicted in Fig. S1(a), which we refer to as the X, Y, and Z modes, respectively, indicating the axis of dominant width oscillations. In the X (Y) mode, the condensate exhibits out-of-phase (in-phase) width oscillations along the $x$ and $y$ directions. We experimentally identified the two quadrupole oscillation modes by applying trap modulations at various frequencies and measuring the width oscillations of the condensate along the $x$ and $y$ directions [Figs. S1(b) and S1(c)]. For our samples at the lowest temperature $\zeta_{\text{th}}<0.1$ and $S=1$, the oscillation frequencies were measured and found to be $\omega/2\pi=7.0(1)$ Hz and $15.0(2)$ Hz for the X and Y modes, respectively, which are in good agreement with the predicted values of $7.0$ Hz and $14.7$ Hz from Eq. (S7) with the trapping frequencies of $(\omega_{x},\omega_{y},\omega_{z})/2\pi=(4.4,8.9,415)$ Hz.
.4 Quantum Reynolds-Averaged Navier-Stokes equation
@@
$-\overline{nv^{\prime}_{\alpha}v^{\prime}_{\beta}}=\overline{n}\nu_{\text{T}}\sigma_{\alpha\beta}-\frac{2}{3}k\delta_{\alpha\beta},$ (S11)
@@
.5 Turbulent damping of collective oscillations
@@
[…]
@@
[…]
To calculate the damping rate of our quadrupole mode, assuming that the amplitude of the quadrupole oscillation does not significantly changed over an oscillation period, the mean energy dissipation rate is expressed as [14]
[…]
where $\langle\ldots\rangle_{T}$ denotes the time average over a period of one oscillation.
@@
with $\bar{R}=(R_{x}R_{y}R_{z})^{1/3}$ and $\bar{\omega}=(\omega_{x}\omega_{y}\omega_{z})^{1/3}$, where $R_{\alpha}$ denotes the TF radius along the $\alpha$ axis. We note that $F(\omega_{\alpha},b_{\alpha})$ is a dimensionless parameter that is determined by the trapping frequencies and the collective oscillation profile, representing the degree of shear stress in the given collective excitation mode.
In Fig. S2, we plot the values $F$ for the X and Y quadrupole modes as functions of the aspect ratio of the transverse trapping frequencies $\lambda=\omega_{y}/\omega_{x}$. In general, the Y mode, which exhibits in-phase oscillation on the $xy$ plane, displays stronger shear stress than the X mode with out-of-phase oscillations, due to the motion along the $z$ direction. This analysis suggests that the dependence of $\Gamma_{\text{T}}$ on the trap geometry and collective modes could provide experimental validation of our turbulent viscosity model for the damped oscillations of a trapped BEC with turbulence.

### v2

@@ Abstract
Finally, we note that the turbulent flow in our system is predominantly two-dimensional due to the oblate trap, whereas the mean flow associated with the collective modes remains three-dimensional. This disparity may limit the use of the standard RANS framework, which presupposes three-dimensional homogeneous isotropic turbulence. Studies of classical turbulence driven by two-dimensional forcing have shown that introducing stronger three-dimensionality can increase the eddy viscosity by suppressing the inverse energy cascade [62]. Mapping how the effective turbulent viscosity evolves across the 2D-3D crossover therefore represents a promising avenue for future work.
@@
In summary, by continuously driving spin dynamics to maintain a steady turbulent state, we have observed an enhancement of damping in the collective oscillations of a turbulent BEC. Viewing this excess through the lens of an effective turbulent viscosity provides a new framework for probing energy-transfer pathways and hydrodynamic properties in superfluid turbulence. This work can extend to other collective excitation modes, particularly including a persistent monopole (breathing) mode available in spherically symmetric BECs [63].
@@ Acknowledgements.
This work was supported by the National Research Foundation of Korea (Grants No. RS-2023-NR077280, No. RS-2023-NR119928, and No. RS-2024-00413957).
Supplemental Material for
Enhancement of damping in a turbulent atomic Bose-Einstein condensate
A Methods
1 Sample preparation
We prepared a thermal cloud of 23Na atoms in the $|F=1,m_{F}=-1\rangle$ state in an optical dipole trap (ODT) and applied evaporative cooling by reducing the trap depth to produce a Bose-Einstein condensate (BEC). The thermal fraction $\zeta_{th}$ of the sample was controlled by adjusting the final value $U_{f}$ of the trap depth. At our lowest temperatures, the typical number of atoms for the BEC was approximately $7\times 10^{6}$. The atomic sample was held in the ODT for several seconds to relax residual motion before driving the collective excitations of the trapped BEC.
Depending on the target experiment, the atomic sample was transformed into a two-component system with $|m_{F}=\pm 1\rangle$ states ($\mathcal{D}_{s}=2$) or a turbulent system with $|m_{F}=0,\pm 1\rangle$ states ($\mathcal{D}_{s}=3$). For the two-component system, we first transferred atoms to the $|m_{F}=0\rangle$ state using a rapid adiabatic passage with an RF magnetic field, followed by a pulse of a strong magnetic field gradient to purify the spin state. We then applied a $\pi/2$ resonant RF magnetic pulse to form an equal mixture of the $m_{F}=\pm 1$ components. After the preparation of the spin mixture, we irradiated a continuous microwave field detuned from the $F=1\rightarrow F=2$ transition to shift the quadratic Zeeman energy to a small negative value. This prevents the creation of atoms in the $|F=1,m_{F}=0\rangle$ state via spin exchange processes, stabilizing the two-component system [1, 2].
The turbulent sample was prepared with the continuous application of the resonant RF field [3]. A steady turbulent state with an irregular spin texture was generated within 0.5 s, exhibiting equal populations of the three $m_{F}=0,\pm 1$ spin components. This turbulence is sustained by chaotic spin dynamics under RF spin driving [4], as detailed in Section B. Following 2 s of RF driving, we conducted the collective oscillation experiment while maintaining the RF spin driving. The BEC lifetime exceeded 30 s under RF driving ($\mathcal{D}_{s}=3$) and was approximately 5 s under the microwave dressing ($\mathcal{D}_{s}=2$), significantly longer than our measurement duration, which lasted for a few hundreds of ms.
@@ 1 Sample preparation
2 Imaging
An absorption image of the sample was taken in the $z$ direction after a time of flight of $\tau=18$ ms by releasing the trapping potential. The number of atoms contained in the thermal gas, $N_{th}$, was estimated from a two-dimensional Gaussian distribution fit to the outer region of the cloud, and the number of condensed atoms $N_{c}$ was determined from an image obtained by subtracting the fitted thermal profile from the original image. The thermal fraction is given by $\zeta_{th}=N_{th}/N$ with $N=N_{c}+N_{th}$. The width $W_{\alpha}(t)$ of the condensates along the $\alpha$ axis ($\alpha=x,y,z$) was determined by fitting the two-dimensional Thomas-Fermi profile to the subtracted image.
3 Analysis of width oscillations
@@
The second term on the right hand side accounts for the width change during the time of flight, owing to the oscillation velocity of the condensate. Here we neglect the expansion effect due to the mean-field energy and quantum pressure, which are negligibly small along the $y$ direction for our highly oblate geometry. When a trapped BEC oscillates as $W_{i}(t)=W_{\text{eq}}(1+B\sin(\omega t+\theta))$, using Eq. (S1), the measured width is given by $W(t)=W_{\text{eq}}(1+\alpha_{\tau}B\sin(\omega t+\theta+\varphi_{\tau}))$ with $\alpha_{\tau}=\sqrt{1+(\omega\tau)^{2}}$ and $\varphi_{\tau}=\tan^{-1}(\omega\tau)$ account for the effects of amplitude amplification and the phase change, respectively, due to time-of-flight expansion. $W_{\text{eq}}$ denotes the in-situ width of the condensate at equilibrium.
In our data analysis, we used the normalized width $\widetilde{W}=W/W_{\text{eq}}$ to account for variations in the number of atoms in the sample. Using the relationship of $W_{\text{eq}}\propto U_{f}^{-1/5}N_{c}^{1/5}$ in the Thomas-Fermi (TF) approximation, the equilibrium width for a given $N_{c}$ was estimated as $W_{\text{eq}}(N_{c})=\overline{W}_{\text{eq}}(N_{c}/\overline{N_{c}})^{1/5}$, where $\overline{W}_{\text{eq}}$ and $\overline{N_{c}}$ are the mean width and the mean number of atoms of the condensate, as determined by averaging tens of measurements without trap modulations.
B Stationary spin-superflow turbulence
@@
$\bm{\nabla}\times\bm{v}=-\frac{i\hbar}{m}\bm{\nabla}\bm{\zeta}^{\dagger}\times\bm{\nabla}\bm{\zeta},$ (S3)
@@
C Quadrupole modes of highly oblate condensates
@@
Figure S1: (a) Quadrupole oscillation modes for an oblate BEC with trapping frequencies $\omega_{x}<\omega_{y}<\omega_{z}$. The oscillatory motion is characterized by $(b_{x},b_{y},b_{z})$, as shown in Eq. (S9), and the three modes are denoted by X, Y, and Z, respectively, corresponding to the dominant oscillation axis. The variation in the density profile in the $xy$ and xzxz planes for each mode is presented. (b),(c) Experimental observation of the quadrupole oscillations of BECs. BEC samples with $\mathcal{D}_{s}=1$ were prepared at our lowest temperatures and the X or Y mode was selectively excited by short trap modulations at a frequency slightly red-detuned from resonance. The upper panels in (b) and (c) display time-of-flight images of BECs oscillating in the X and Y modes, respectively, at different hold times. The lower panels show the time evolution of the normalized condensate widths in the $x$ and $y$ directions, respectively. Data points represent single measurements and the solid lines indicate damped sinusoidal fits to the experimental data.
@@
The three quadrupole excitation modes for an oblate trapping potential with $\omega_{x}<\omega_{y}<\omega_{z}$ are depicted in Fig. S1(a), which we refer to as the X, Y, and Z modes, respectively, indicating the axis of dominant width oscillations. In the X (Y) mode, the condensate exhibits out-of-phase (in-phase) width oscillations along the $x$ and $y$ directions. We experimentally identified the two quadrupole oscillation modes by applying trap modulations at various frequencies and measuring the width oscillations of the condensate along the $x$ and $y$ directions [Figs. S1(b) and S1(c)]. For our samples at the lowest temperature $\zeta_{th}<0.1$ and $\mathcal{D}_{s}=1$, the oscillation frequencies were measured and found to be $\omega/2\pi=7.0(1)$ Hz and $15.0(2)$ Hz for the X and Y modes, respectively, which are in good agreement with the predicted values of $7.0$ Hz and $14.7$ Hz from Eq. (S7) with the trapping frequencies of $(\omega_{x},\omega_{y},\omega_{z})/2\pi=(4.4,8.9,415)$ Hz.
D Resonance behavior of quadrupole oscillations
[…]
We investigated the resonance of the Y quadrupole mode with trap modulations of $U(t)=U_{f}[1+\epsilon\sin(\omega_{d}t)]$ by measuring the steady-state response of the BEC as a function of the modulation frequency $\omega_{d}$. In this case, we applied the trap modulation for a time of $t_{0}=50\frac{2\pi}{\omega_{d}}$, ensuring that the BEC oscillations reach a quasi-steady state, and continued the modulation in the subsequent hold time $t_{h}$. We tracked the normalized width $\widetilde{W}(t_{h})$ of the BEC along the $y$ direction by taking a time-of-flight image of the sample for various modulation times $t_{h}$ [Fig. S2(a)], and determined the in-trap relative amplitude $B$ and phase $\theta$ of the BEC oscillations (see Section 3).
The resonance behavior of the driven BEC was characterized with the amplitude magnification factor $\widetilde{A}(\omega_{d})$, which is the ratio of the oscillation amplitude $B$ to the equilibrium displacement for the applied forcing. Since the change of $W_{\text{eq}}(\propto U_{f}^{-1/5})$ for small $\Delta U=\epsilon U_{f}$ is $\Delta W_{\text{eq}}=-\frac{1}{5}\epsilon W_{\text{eq}}$, the factor is estimated as $\widetilde{A}=5B/\epsilon$. To ensure our measurements were in the linear response regime, we kept the oscillation amplitude $B$ below 5%. We verified that the oscillation amplitude $B$ increases linearly with $\epsilon$ at a modulation frequency of $\omega_{d}/2\pi=27$ Hz in this limit.
Figure S2 shows the response spectra $\widetilde{A}(\omega_{d})$ and $\theta(\omega_{d})$ for three different samples, including non-turbulent BECs prepared without RF spin driving, with respective thermal fractions of $\zeta_{th}=0.32(1)$ (blue markers) and 0.59(1) (cyan), and a turbulent BEC with $\zeta_{th}=0.57(1)$ (orange). The trapping frequencies were $(\omega_{x},\omega_{y},\omega_{z})=2\pi\times(7.5,15.3,730)$ Hz, giving the quadrupole mode frequency $\omega_{\nu,0}/2\pi=25.3$ Hz at zero temperature. The resonance behavior is evident in $\widetilde{A}$, accompanied by a phase change of $\pi$ in $\theta$ as $\omega_{d}$ increases. The resonance frequency $\omega_{\nu}$ and the spectral width $\Gamma_{s}$ are determined from the simultaneous fit of $\widetilde{A}(\omega_{d})=\frac{A^{\prime}\omega_{\nu}^{2}}{\sqrt{(\omega_{d}^{2}-\omega_{\nu}^{2})^{2}+4(\omega_{d}\Gamma_{s})^{2}}}$ and $\theta(\omega_{d})=\frac{\pi}{2}+\tan^{-1}(\frac{\omega_{d}^{2}-\omega_{\nu}^{2}}{2\Gamma_{s}\omega_{d}})-\theta_{0}$ to the experimental data with $A^{\prime}$ and $\theta_{0}$ being free parameters. In the fitting, $A^{\prime}\approx 1.7$ and $\theta_{0}\approx 0.1\pi$. The small non-zero value of $\theta_{0}$ is attributed to a systematic effect in the time-of-flight measurement. For BECs without turbulence, we obtain $\Gamma_{s}/2\pi=0.40(3)$ Hz and $\Delta\omega_{\nu}/2\pi=(\omega_{\nu}-\omega_{\nu,0})/2\pi=0.21(3)$ Hz for a low $\zeta_{th}$, and $\Gamma_{s}/2\pi=0.53(5)$ Hz and $\Delta\omega_{\nu}/2\pi=0.44(1)$ Hz for a high $\zeta_{th}$. As $\zeta_{th}$ increases, the resonance frequency shifts upward and the damping increases. The turbulent BEC demonstrates a more pronounced spectral broadening and a larger shift with $\Gamma_{s}/2\pi=1.1(1)$ Hz and $\Delta\omega_{\nu}/2\pi=1.6(3)$ Hz.
We briefly comment on the frequency shifts observed in the driven response spectra. Anomalous frequency shifts and damping of collective modes were first reported in Ref. [11], and subsequent studies identified mean-field coupling between the condensate and thermal cloud as the primary mechanism [12, 13]. In particular, the direction of the frequency shift can reverse depending on their relative motion [12]. In our driven experiments, the observed blueshift is probably due to this coupling, indicating that a simple one-way Landau damping picture is insufficient. A full explanation would require modeling the long-term dynamical interaction between the condensate and the thermal cloud, such as within the Zaremba-Nikuni-Griffin (ZNG) formalism [12], which is beyond the scope of this work. We note that in our free-decay experiments, where a short pulse drive is applied at an optimized frequency, no significant temperature-dependent frequency shift was observed.
In Fig. S3, we present the measured values of $\Gamma_{s}/\omega_{\nu}$, together with the relative damping rates $\Gamma/\omega_{\nu}$ from Fig. 3(b). The slightly higher values of $\Gamma_{s}/\omega_{\nu}$ are probably due to the steady-state relative motion between the condensate and the thermal cloud during the driven measurements.
Figure S3: Relative spectral width $\Gamma_{s}/\omega_{\nu}$ (open markers) as a function of reduced temperature $\widetilde{T}$, extracted from the response spectra shown in Fig. S2. For comparison, the solid circles represents the relative damping rates $\Gamma/\omega_{\nu}$ from free-decay measurements, as also shown in Fig. 3(b). The solid and dashed lines are the same as those described in the caption of Fig. 3.
E Quantum Reynolds-Averaged Navier-Stokes equation
@@
$\overline{nv^{\prime}_{\alpha}v^{\prime}_{\beta}}=\overline{n}\nu_{\text{T}}\sigma_{\alpha\beta}-\frac{2}{3}k\delta_{\alpha\beta},$ (S11)
@@
F Turbulent damping of collective oscillations
@@
The primary effect of turbulent viscosity is the dissipation of energy from the quadrupole oscillations. The total mechanical energy $E$ for the collective oscillations with respect to the equilibrium state is twice the time average of the kinetic energy $E_{kin}(t)=\int d\bm{r}\frac{1}{2}m\overline{n}(\bm{r},t)\overline{v}^{2}(\bm{r},t)$ [16] and up to the linear order of $\delta\overline{n}$ and $\overline{\bm{v}}$, it is given by
$E=\int d\bm{r}\left[\frac{1}{2}mn_{0}(\bm{r})\sum_{\alpha}\left(\frac{2g}{m\omega}\right)^{2}b_{\alpha}^{2}r_{\alpha}^{2}\right].$ (S12)
@@
$\dot{E}(t)=\int d\bm{r}\sum_{\alpha,\beta}m{\overline{v}_{\alpha}}\frac{\partial}{\partial r_{\beta}}\left(\overline{n}\nu_{\text{T}}\sigma_{\alpha\beta}-\frac{2}{3}k\delta_{\alpha\beta}\right).$ (S13)
To calculate the damping rate of our quadrupole mode, assuming that the amplitude of the quadrupole oscillation does not significantly changed over an oscillation period, the mean energy dissipation rate is expressed as
$\langle\dot{E}\rangle_{T}=-\int d\bm{r}\left[\frac{1}{2}mn_{0}(\bm{r})\nu_{\text{T}}\sum_{\alpha,\beta=1}^{3}\left\langle\sigma_{\alpha\beta}^{2}(\bm{r},t)\right\rangle_{T}\right],$ (S14)
where $\langle\ldots\rangle_{T}$ denotes the time average over a period of one oscillation [17].
@@
with $\bar{R}=(R_{x}R_{y}R_{z})^{1/3}$ and $\bar{\omega}=(\omega_{x}\omega_{y}\omega_{z})^{1/3}$, where $R_{\alpha}$ denotes the TF radius along the $\alpha$ axis. $F(\omega_{\alpha},b_{\alpha})$ is a dimensionless parameter that is determined by the trapping frequencies and the collective oscillation profile, representing the degree of shear strain in the given collective excitation mode.
In Fig. S4, we plot the values $F$ for the X and Y quadrupole modes as functions of the aspect ratio of the transverse trapping frequencies $\lambda=\omega_{y}/\omega_{x}$. In general, the Y mode, which exhibits in-phase oscillation on the $xy$ plane, displays stronger shear stress than the X mode with out-of-phase oscillations, due to the motion along the $z$ direction [Fig. S1(a)]. For the trap parameters used in our experiment, we find $F=0.345$ for the Y mode and $F=0.096$ for the X mode, implying that $\Gamma_{\text{T}}$ for the Y mode is approximately three times greater than that of the X mode for a given $\nu_{\text{T}}$.
For comparison, we also measured the damping rate of the X mode. The experimental protocol was identical to that used for the Y mode, except that the drive frequency $\omega_{d}$ was tuned specifically for the X mode, and the condensate width $\widetilde{W}$ was measured along the $x$ direction, where the X mode exhibits a larger oscillation amplitude [Fig. S1(b)]. The trap frequencies were $(\omega_{x},\omega_{y},\omega_{z})=2\pi\times(7.5,15.3,730)$ Hz. For $\mathcal{D}_{s}=3$ turbulent samples at a reduced temperature of $\widetilde{T}=4.6(1)\times 10^{-3}$, we measured $\Gamma/2\pi=0.48(11)$ Hz and $\omega_{\nu}/2\pi=11.7(1)$ Hz. For $\mathcal{D}_{s}=1$ samples at $\widetilde{T}=6.0(2)\times 10^{-3}$, the corresponding values were $\Gamma/2\pi=0.21(7)$ Hz and $\omega_{\nu}/2\pi=11.6(1)$ Hz.
Using the relation $A_{\nu}^{(0)}=\Gamma/(\omega_{\nu}\widetilde{T}) fo$r the $\mathcal{D}_{s}=1$ case, the excess damping in the $\mathcal{D}_{s}=3$ turbulent sample is determined from $\Gamma_{\text{T}}=\Gamma-3A_{\nu}^{(0)}\widetilde{T}\omega_{\nu}$ [Eq. (2)] as described in the main text. From the two measurements, we extract the damping rate of $\Gamma_{\text{T}}=-0.01\pm 0.19$ Hz. Given that $\Gamma_{\text{T}}/2\pi\approx 0.2$ Hz was observed for the Y mode, we expect $\Gamma_{\text{T}}/2\pi\approx 0.06$ Hz for the X mode, which lies below the resolution of our current experiment. A more systematic study exploring the dependence of $\Gamma_{\text{T}}$ on trap geometry and shear stress would be needed to rigorously validate the model.